OpenStudy (anonymous):

How do I find the turning point of y = 6x^2 - 24x + 18 So far I only factored y = 6(x - 1)(x - 3) and dont know where to go from here. @Mathematics

OpenStudy (anonymous):

Anyone?

OpenStudy (saifoo.khan):

Turning point = vertex. \[6x^2-24x+18=0\]can also be written as, \[6(x-2)^2-6\] So, t.p. = (2,-6)

OpenStudy (anonymous):

Oh okay, so I can't factor it, I have to complete the square again. Thanks very much!